OpenStudy (anonymous):

x 25 35 45 55 y 92 86 80 74 What equation that represents this relationship

OpenStudy (greencat):

easy

OpenStudy (greencat):

x increases 10

OpenStudy (greencat):

y decreases 6

OpenStudy (greencat):

slop = rizse/run

OpenStudy (greencat):

slope*

OpenStudy (greencat):

slop=-6/10

OpenStudy (greencat):

slope=-0.6

OpenStudy (greencat):

y=-0.6x+b

OpenStudy (greencat):

plug in

OpenStudy (greencat):

80=-0.6(45)+b

OpenStudy (greencat):

80=-27+b

OpenStudy (greencat):

b=80+27

OpenStudy (anonymous):

thank you :)

OpenStudy (greencat):

b=107

OpenStudy (greencat):

y=mx+b

OpenStudy (greencat):

y=-.06x+90

OpenStudy (greencat):

ur welcome

OpenStudy (greencat):

thanks

OpenStudy (anonymous):

You're welcome. :>